30 60 to 0,05F t) 12V 2ix Determine v for tzo.
Answer #1

31 0.05FT DIE 12V VCD = ? veete VI) - Vm)-[vvetlo 120 5-0 (steady state] Capacitos – Oper ckt VLO) VOJVCO)= 12v 676 6 Br 12vE 200 (steaty State) VCO) 0 V(Co)-2 Apply kvL los dos 12-2x +biga x=-34) (Val- T=Resc Reas las 202 (21 E Regg- Res Voc Voc - vol. or Voc - Vol. v + kve bar top -12-61x +3 Is=0 3 Isc = 2ix – sie speltage erval -12-bix +2 = Ise a 2x (2=– T-Reac = 3 × 0.05 Tois sed Vethe VC)-[urce) -vcem --6-5-6-18) et TVA)=-6 +180 tot zo